OpenStudy (anonymous):

verify (cosx+sinx)/(cosx-sinx)=sec2x+tan2x

OpenStudy (anonymous):

please!

OpenStudy (anonymous):

this didn't fall out like I thought it was gonna, sorry. 1 sec.... Above though, since sec^2 = 1/cos^2 and tan^2 = sin^2 / soc^2 you can rewrite it like that

OpenStudy (anonymous):

but, its not sec^2 or tan^2x its sec2x and tan2x, they are double angles

OpenStudy (anonymous):

well then ignore me completely :P

OpenStudy (anonymous):

haha alright. Can you still help me?

OpenStudy (anonymous):

yeah, it might make it easier. 1 sec \[ \frac{\cos x + \sin x}{\cos x - \sin x} = \frac{1+\sin 2x}{\cos 2x}\]

OpenStudy (anonymous):

could you explain how you got 1+sin2x/cos2x?

OpenStudy (anonymous):

yah \[ \sec2 x = \frac{1}{\cos2 x} ; \ \ \tan2x = \frac{\sin2x}{\cos2x}\] they have a common denominator

OpenStudy (anonymous):

phew, ok

OpenStudy (anonymous):

you'll need double angle identities \[ \sin 2x = 2 \sin x \cos x \\ \cos 2x = \cos^2 x - \sin^2 x\] Then cross multiply \[ \cos 2x ( \cos x + \sin x) = (1+ \sin 2x)(\cos x - \sin x)\] and simplify. 6 lines or so. You should end up with 1=1 Have a go at it and see if you get caught up anywhere ^_^

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

^_^